Calculate the enthalpy of formation of ethanol (C2H5OH) given the following enthalpies of combustion.
ΔcH C(s) = -393, H2(g) = -286, C2H5OH(l) = -1371 kJ/mol
1
Expert's answer
2020-05-07T14:16:02-0400

"4C+6H_2+O_2\to 2C_2H_5OH" "\Delta fH=?"

"C+O_2\to CO_2" "\Delta cH(C)=-393kJ/mol"

"2H_2+O_2\to2H_2O" "\Delta cH (H_2)=-286 kJ/mol"

"C_2H_5OH+3O_2\to 2CO_2+3H_2O" "\Delta cH(C_2H_5OH)=-1371kJ/mol"

"\Delta fH=4\Delta cH(C)+6\Delta cH(H_2)-2\Delta cH(C_2H_5OH)"

"\Delta fH=" 4*(-393kJ/mol)+6*(-286kJ/mol)-2(-1371kJ/mol)=-546kJ/mol

Answer:-546kJ/mol